Draw the parabola with the equation: 4x

2 + 6x − 3y = 4.
1
Expert's answer
2020-10-14T18:10:45-0400

Given equation is

"4x^2+6x-3y=4"


"4(x^2+\frac{3}{2}x)-3y=4"


"4(x^2+\frac{3}{2}x+\frac{9}{16}-\frac{9}{16})-3y=4"


"4(x+\frac{3}{4})^2-\frac{9}{4}-3y=4"


"4(x+\frac{3}{4})^2=3y+4+\frac{9}{4}"


"4(x+\frac{3}{4})^2=3y+\frac{25}{4}"


at x=0, "4\times \frac{9}{16}=3y+\frac{25}{4}"


"3y=\frac{9}{4}-\frac{25}{4}"


"3y=\frac{-16}{4}"


3y=-4

y=-1.33


At y=0,

"4(x+\frac{3}{4})^2=\frac{25}{4}" ,"(x+\frac{3}{4})^2=\frac{25}{16}" ,"(x+\frac{3}{4})=\pm\frac{5}{4}"


"we get x=\frac{1}{2}" and "-2"

our parabola must pass through (0.5,0),(-2,0) and (0,-1.33)

The parabola is